OpenStudy (anonymous):

find the volume of the solid bounded by z=-y, the xy plane, and r=4cos(theta)

terenzreignz (terenzreignz):

We need a function, it seems best to use polar coordinates no?

OpenStudy (anonymous):

yes

terenzreignz (terenzreignz):

So, convert z = -y to \[\huge z = -r\sin \theta\]

OpenStudy (anonymous):

and xy plane mean lower limit of z =0

terenzreignz (terenzreignz):

Wait, no.

terenzreignz (terenzreignz):

The function to integrate is z... \[\huge z = f(r,\theta) = -4\sin \theta\]

terenzreignz (terenzreignz):

Wait... stop... we'll redo this.

terenzreignz (terenzreignz):

(sorry, I actually just woke up :) )

terenzreignz (terenzreignz):

Let's do it in rectangular instead...

terenzreignz (terenzreignz):

And convert \[\large r = 4 \cos \theta\]into rectangular.

terenzreignz (terenzreignz):

Is it even possible?

OpenStudy (anonymous):

you mean we convet it into rectangular respect to x,y,z?

terenzreignz (terenzreignz):

Yeah...

OpenStudy (anonymous):

or cylindrical r, theta and z?

terenzreignz (terenzreignz):

wait, let me have a good think...

OpenStudy (anonymous):

ok, take your time, I appreciate your help

terenzreignz (terenzreignz):

Okay, definitely better to use rectangular. But first, I need you (not really, lol, but do it anyway :) ) to convert \[\huge r = 4\cos \theta\] into rectangular form.

OpenStudy (anonymous):

x = r costheta= 4cos^2 theta?

terenzreignz (terenzreignz):

yeah... using those, and also \[\huge r = \sqrt{x^2 + y^2}\]

OpenStudy (anonymous):

y =4cos sin

terenzreignz (terenzreignz):

uhh... be a little more clever :P \[\huge \cos \theta = \frac{x}{r}\]

OpenStudy (anonymous):

don't get, hihihi the purpose of this step

terenzreignz (terenzreignz):

Well... \[\huge r = 4\left(\frac{x}{r}\right)\] \[\huge r^2 = 4x\] \[\huge x^2 + y^2 = 4x\]

OpenStudy (anonymous):

oh, to get z = x^2 +y^2 -4x?

terenzreignz (terenzreignz):

No... this is a region, \[\huge x^2 - 4x + y^2 = 0\]

terenzreignz (terenzreignz):

\[\huge (x-2)^2 + y^2=4\]

OpenStudy (anonymous):

yes, but it is a upper limit of z, too?

terenzreignz (terenzreignz):

This is just the region over which we integrate, the real function that we're integrating is \[\huge z = f(x,y) = -y\]

OpenStudy (anonymous):

ok, you have a circle there

terenzreignz (terenzreignz):

yes.

terenzreignz (terenzreignz):

\[\huge \iint\limits_R -y \ dA\]

OpenStudy (anonymous):

so the function to take integral yes, got it ,

terenzreignz (terenzreignz):

Where R is the region bounded by that circle.

OpenStudy (anonymous):

tripple or double?

terenzreignz (terenzreignz):

Just double. We're only taking a volume.

OpenStudy (anonymous):

what different between double and tripple, when we use this or that? both them compute volume ?

terenzreignz (terenzreignz):

Triple Integral actually computes Hyper-volume, but may be used to compute volume. We can solve this using a triple integral, but it will lead to this double integral anyway.

OpenStudy (anonymous):

wow, just get from you

terenzreignz (terenzreignz):

hmm?

OpenStudy (anonymous):

we just have the function, how to set the limit?

terenzreignz (terenzreignz):

We have the region... R the circle defined by \[\huge (x-2)^2+y^2 = 4\]

OpenStudy (anonymous):

do we have to convert back to r and theta?

terenzreignz (terenzreignz):

No. You can forget about those, actually :)

OpenStudy (anonymous):

why? it's easier, right? we have circle r from 0 to2, theta from 0 to 2pi and the function is -y means -rsin theta, is it not enough to compute?

terenzreignz (terenzreignz):

Yes, unfortunately, the usefulness of using polar coordinates diminishes rapidly when your circle is not centred around the pole (origin)

OpenStudy (anonymous):

really? oh yea, you are right!!! :)

terenzreignz (terenzreignz):

I told you... it's Terence :D Now, the region has y running from \(\large -\sqrt{4-(x-2)^2}\) to \(\large \sqrt{4-(x-2)^2}\)

OpenStudy (anonymous):

hihihi.... yes, Terence!!!

terenzreignz (terenzreignz):

And x runs from 0 to 4

terenzreignz (terenzreignz):

Because the circle is centred around the point (2,0)

OpenStudy (anonymous):

got it.

OpenStudy (anonymous):

so the whole thing is int_from 0 to 4 int_ (the bunch you give out above) of -y dy dx

terenzreignz (terenzreignz):

\[\huge \int_{0}^4\int_{-\sqrt{4-(x-2)^2}}^{\sqrt{4-(x-2)^2}}-y \ dy dx\]

terenzreignz (terenzreignz):

Might look hard, but this integral should be easy...

OpenStudy (anonymous):

hi Terence, how can you have that keen flexibility, how can you figure out right after look at the problem and decide which method should be use, teeeaaach me , please

terenzreignz (terenzreignz):

A diet of noodles? I really don't know... it sort of just happens :) I'd say experience, but I don't have that much. I do stand by that epithet, though... experience is the best teacher :D

OpenStudy (anonymous):

Lol, got it, thanks a lot Terence,

terenzreignz (terenzreignz):

No problem :) I'm going to sign off for now. --------------------------------------------- Terence out